OpenStudy (anonymous):

Write the integral that produces the same value as the limit as n goes to infinity of the summation from i equals 1 to n of the product of the quantity 1 plus 3 times i over n and 3 over n.

OpenStudy (green_1):

@RachelDude @Liam_Monge @lakota1999

OpenStudy (anonymous):

Is this your sum?\[\lim_{n\to\infty}\sum_{i=1}^n \frac{1+3i}{n}\times\frac{3}{n}\]